Can someone please help me with these. A.S.A.P.

1) Indicate the formula for the following conditions.
P(X | Y) = ________
2) Type the general term for the following sequence.
3, 9, 27, 81, 243, . . . __________
3) Type the general term for the following sequence.
3/4, 6/5, 9/6, 12/7, 15/8 ...
4) Evaluate the following expression.
Pc(5,5) = ______
5) A person is wearing a bracelet with 11 settings around the bracelet. How many different ways can 11 birthstones be arranged around the bracelet? Leave the answer in factorial form. ________
6) How many different ways can an executive committee of nine people sit at a round table? _________
7) A compound experiment consists of drawing at random twice without replacing from a jar with 3 red, 3 white, and 2 black marbles. What is the probability of getting first a red and then a black marble? ________

Thank You,
~Sarah :)

Answer by rfer(16322)

Here is part of it for you.
2)Start with 3 and multuply by 3 repeatedly.
3)Numerator: Start with 3 and add 3 repeatedly
Denominator: Start with 4 and add 1 repeatedly.
6) (9)(9)=81 ways
7) (3/8)(2/7)=6/56=3/28